EDITORIAL

Terry Goodman

We are busy here at MJMS as we attend to all the details of completing another semester as well as putting this issue of the journal together. As mentioned in a previous editorial, we will be implementing a number of changes over the next year or so. We anticipate that these changes will improve MJMS and help us to better serve our readers. We continue to solicit your input concerning these proposed changes.

Over the years we have attempted to publish articles that focus on a wide variety of mathematics topics and areas of study. At the present time, we are particularly interested in receiving manuscripts that focus on topics in geometry, computer science, mathematics education, and applied mathematics. We also want to encourage you to continue to contribute to our Problems sections; both problems and solutions. This has been one of the most popular features of the journal and we look forward to future problems and solutions.

This fall, the Department of Mathematics and Computer Science at CMSU in cooperation with the Education Satellite Network, will be broadcasting, via satellite, a History of Mathematics course. The course will target middle grade mathematics teachers as its audience, but we anticipate that many high school mathematics teachers will also participate. This course will be taught by Robert Kennedy (Coordinating Editor for MJMS). We anticipate that the course will be broadcast to 20 Host School Sites around the state of Missouri. Other locations (not limited to Missouri) will be able to tune in to the broadcasts if they have the appropriate downlink equipment and knowledge of the satellite settings and coordinates. If you would like to have this information, please contact me by phone, mail, or e–mail. MJMS would also welcome manuscripts that focus on the historical development of mathematics and/or emphasize ways to incorporate ideas from the history of mathematics and that can be used to help students see and appreciate the nature, role, and fascination of mathematics.

If you know of upcoming special courses, workshops, seminars, or conferences in your area we would encourage you to share such information with us. We will be pleased to make this information available to our readers.